# Predicting Change, Not States: An Alternate Framework for Neural PDE Surrogates

### v1

@@ 5.1 Loss Landscapes
To understand why derivative prediction may outperform state prediction, we train a Unet on the 1D Advection benchmark using both state and derivative prediction, and visualize the loss landscape using filter normalization [52] in Figure 5. The loss landscape is constructed by perturbing trained weights in a set of normalized, random directions and evaluating the perturbed rollout loss on validation samples. We find that the derivative prediction objective can be empirically easier to learn than the state prediction objective. One potential reason for this is that when solving PDEs at a suitable resolution, future timesteps are often similar to current ones but with slight changes from evolving forward in time. This causes the prediction label to contain redundant information about the current state, where the learning signal comes from the change in state, which can be small. This can contribute to many local minima, where models can learn variations of the identity map that predict potential states that resemble the ground truth, since the forward evolution only slightly changes the solution [27]. However, despite nodal values being similar among local minima, the temporal derivatives at these potential states can be very different. Therefore, derivative prediction can isolate the important dynamics to better distinguish between potential model predictions.
@@
Figure 6: Snapshots of the 2D Navier-Stokes equation, along with their temporal derivatives. Adding noise to a trajectory creates a similar sample, however this effect is noticeable when taking its temporal derivative.
Figure 7: Error at each timestep using FNO, Unet, or an oracle to estimate $\frac{\partial\mathbf{u}}{\partial t}$ and integrating with the Forward Euler method, for the 1D Advection and 2D NS equations. Each marker denotes an integration step.
5.2 Numerical Error
We have previously shown that numerical error can cause models trained with derivative prediction to become unstable at large $\Delta t$, however, we seek to further understand this behavior. In particular, we have access to ground-truth derivative estimates from the ground-truth trajectory; using these estimates rather than neural surrogate estimates can serve as a baseline to compare how much error is being contributed by the ODE integrator. This baseline is called the numerical oracle, since it relies on future observations to accurately estimate the instantaneous derivative.
@@ 5.2 Numerical Error
From training, we know that the one-step training loss is low and can assume that if given a perfect input, models can make an accurate derivative estimate. Indeed, this seems to be the case where the model errors closely follow the oracle errors up to a certain point (around the first $20\%$ of the rollout). However, after this point, the numerical error shifts model inputs such that they no longer match the training distribution, and the model begins to make inaccurate derivative estimates, which contributes model error. Integrating these inaccurate estimates produces model inputs even further from the training distribution, which produces even more inaccurate derivative estimates. This error compounds until the model no longer makes useful predictions.
@@ 6 Conclusion
We introduce framing neural surrogates to predict temporal derivatives and use an ODE integrator during inference to produce solutions. Although this idea has been studied informally through residual prediction, this is the first work to formally consider this framework with an ODE integrator and investigate its benefits and limitations. We find that derivative prediction can improve model accuracy and stability across architectures and PDE problems, and has useful data and inference flexibility. Furthermore, we compare this framework with other training modifications and study the effect of temporal resolution on model performance. Lastly, we hypothesize potential explanations for the improved stability and accuracy of derivative prediction and study the role of numerical error introduced by the ODE integrator. We hope that future work can apply derivative prediction across a wide variety problems and model architectures, and improve performance at coarser step sizes.

### v2

@@ 5.1 Loss Landscapes
To understand why derivative prediction may outperform state prediction, we train a Unet on the 1D Advection benchmark using both state and derivative prediction, and visualize the loss landscape using filter normalization [74] in Figure 5. The loss landscape is constructed by perturbing trained weights along a linear combination of normalized, random directions and evaluating the perturbed rollout loss on validation samples. Specifically, given $\theta^{*}\in\mathbb{R}^{n}$, which are the trained model parameters (n is the parameter count), we define two random vectors $\delta\in\mathbb{R}^{n}$ and $\gamma\in\mathbb{R}^{n}$. Given a loss function $\mathcal{L}_{\theta}(x,y)$, multiple rollout loss values are calculated after varying the network’s parameters $\mathcal{L}_{\theta^{*}+a\delta+b\gamma}(x,y)$ by a linear combination of $\delta$ and $\gamma$. Each time a loss is calculated with a specific set of parameters $\theta^{*}+a\delta+b\gamma$, it is averaged over the validation set $(x,y)\in D_{val}$. We sample $a$ and $b$ along a grid, which form the two axes, with $\mathcal{L}_{\theta^{*}+a\delta+b\gamma}(x,y)$ being the third, vertical axis or contour set.
Figure 6: Noised Trajectories. Snapshots of the 2D Navier-Stokes equation, along with their temporal derivatives. Adding noise to a trajectory creates a similar sample, however this effect is noticeable when taking its temporal derivative.
We find that the derivative prediction objective can be empirically easier to learn than the state prediction objective. One potential reason for this is that when solving PDEs at a suitable resolution, future timesteps are often similar to current ones but with slight changes from evolving forward in time. This causes the prediction label to contain redundant information about the current state, where the learning signal comes from the change in state, which can be small. This can contribute to many local minima, where models can learn variations of the identity map that predict potential states that resemble the ground truth, since the forward evolution only slightly changes the solution [32]. However, despite nodal values being similar among local minima, the temporal derivatives at these potential states can be very different. Therefore, derivative prediction can isolate the important dynamics to better distinguish between potential model predictions.
@@
5.2 Error Analysis
After investigating the performance of different training frameworks across models, PDEs, training modifications, and timescales, we are also interested in examining the sources of error to better understand the mechanisms behind derivative prediction.
Model Adv Heat KS Burgers NS Kolm. Flow
FNO-State (Next-Step Error) 0.0048 0.0053 0.0012 0.0382 0.0016 0.0214
FNO-Derivative (Next-Step Error) 0.0004 0.0020 0.0003 0.0045 0.0005 0.0038
FNO-Derivative (Derivative Error) 0.0204 0.2665 0.0047 0.1104 0.0234 0.0522
Unet-State (Next-Step Error) 0.0041 0.0044 0.0013 0.0291 0.0018 0.0254
Unet-Derivative (Next-Step Error) 0.0003 0.0018 0.0008 0.0068 0.0008 0.0094
Unet-Derivative (Derivative Error) 0.0101 0.3065 0.0142 0.1700 0.0402 0.1256
Table 4: Next-Step Error. Results on validation next-step error across different PDEs, models, and training/inference frameworks. Derivative error is calculated with an RK4 integrator. Next-step error for state-prediction models is given by: $\mathcal{L}(\mathbf{u}(t_{n+1}),F_{\theta}(\mathbf{u}(t_{n})))$, while for derivative-prediction models both the derivative error: $\mathcal{L}(\frac{d\mathbf{u}}{dt}|_{t=t_{n}},F_{\theta}(\mathbf{u}(t_{n})))$ and next-step error: $\mathcal{L}(\mathbf{u}(t_{n+1}),\int F_{\theta}(\mathbf{u}(t_{n})))$ are given.
[…]
Next-Step Error While rollout error and correlation time are presented in the main results, examining next-step error can give information into how well the model is able to learn its training objective. While the transition from next-step error to rollout error is straightforward in state-prediction (rollout error is the sum of next-step error along a trajectory), it is worthwhile to consider the next-step error for derivative-prediction. Specifically, we are interested in how well models can match empirical derivatives and if integrating the learned derivative gives a similar next-step error to state-prediction frameworks. We present these results in Table 4 for all models and PDEs. Next-step error is reported for state-prediction models and derivative-prediction models, averaged over all timesteps in the validation set. For state prediction models, next-step error is given by: $\mathcal{L}(\mathbf{u}(t_{n+1}),F_{\theta}(\mathbf{u}(t_{n})))$, while for derivative-prediction models we report both derivative error: $\mathcal{L}(\frac{d\mathbf{u}}{dt}|_{t=t_{n}},F_{\theta}(\mathbf{u}(t_{n})))$ and next-step error: $\mathcal{L}(\mathbf{u}(t_{n+1}),\int F_{\theta}(\mathbf{u}(t_{n})))$ after integrating the predicted derivative. Furthermore, to compare the losses between the datasets and the models, all errors are calculated using a relative L2 loss, as in Section 4.1.
After decoupling the derivative prediction and ODE integrator, we find that the derivative error is consistently higher than next-step error, yet it is still low enough to achieve better performance than state prediction models. We hypothesize this is because temporal derivatives can be more challenging to predict than solution fields; for example, in Burgers equation, nonlinear terms can cause sharp discontinuities in $\frac{d\mathbf{u}}{dt}$. However, using a sufficiently fine temporal resolution and a higher-order integrator can account for this error by scaling the derivative by $\Delta t$ and using multiple derivative predictions. Both the model and time-stepping scheme can contribute to lower next-step error, and subsequently lower rollout error, for derivative prediction models.
[…]
Numerical Error While models can learn accurate derivative estimates, this is not enough to maintain accuracy over a rollout. In particular, even with access to perfect derivatives, numerical integrators can still accumulate discretization error. Therefore, we seek to understand to what extent this affects derivative prediction rollouts. In particular, we have access to ground-truth derivative estimates from the ground-truth trajectory; using these estimates rather than neural surrogate estimates can serve as a baseline to compare how much error is being contributed by the ODE integrator. This baseline is called the numerical oracle, since it relies on future states to accurately estimate the instantaneous derivative (i.e., when given an initial state, we look into the future to compute $\frac{d\mathbf{u}}{dt}$).
@@ 5.2 Error Analysis
Figure 7: Numerical Error. Error at each timestep using FNO, Unet, or an oracle to estimate $\frac{\partial\mathbf{u}}{\partial t}$ and integrating with the Forward Euler method, for the 1D Advection and 2D NS equations. Each marker denotes an integration step.
From the previous analysis, we know that the next-step validation loss is low and can assume that if given a perfect input, models can make an accurate derivative estimate. Indeed, this seems to be the case where the model errors closely follow the oracle errors up to a certain point (around the first $20\%$ of the rollout). However, after this point, the numerical error shifts model inputs such that they no longer match the training distribution, and the model begins to make inaccurate derivative estimates, which contributes model error. Integrating these inaccurate estimates produces model inputs even further from the training distribution, which produces even more inaccurate derivative estimates. This error compounds until the model no longer makes useful predictions.
@@
5.3 Timing Experiments
A final consideration for derivative prediction is that it can take longer during inference, since higher-order integrators can require multiple predictions per step. Therefore, we compare the computational cost of models using derivative prediction to those using state prediction, as well as to baseline numerical solvers. The results are given in Table 5 for the Navier-Stokes and Kolmogorov Flow experiment. Models are run during inference to predict a full rollout for each validation sample, and times are averaged across the validation set. Furthermore, model sizes are kept consistent across all timing experiments. We use the pseudospectral solver for the Navier-Stokes experiment proposed by Li et al. [3] and a JAX-implemented, pseudospectral ETDRK4 solver for the Kolmogorov Flow experiments, proposed by Koehler et al. [72]. Batch sizes are set to maximize GPU usage across all experiments. Following McGreivy and Hakim [26], we also coarsen the solver resolution to match the neural surrogate accuracy when needed to provide a fair comparison of computational cost. The rollout accuracy of the coarse solution is calculated by using bilinear interpolation to map the coarse solution onto the finer grid, and the relative L2 error is calculated for the interpolated trajectory and averaged across the validation set. All computations are performed on a single NVIDIA RTX 2080 Ti GPU for the Navier-Stokes experiment and a single NVIDIA RTX 6000 Ada GPU for the Kolmogorov Flow experiment.
We find that with lower-order schemes, the computational cost of derivative prediction is not meaningfully higher than state prediction. This is because in the Forward Euler scheme, the additional overhead is extremely small, as we only need to add and scale the predicted derivative to the current value. Likewise, for the Adams-Bashforth scheme, despite using multiple model predictions and having a lower truncation error, the previous predictions are cached and therefore do not need to be re-evaluated, resulting in minimal overhead. However, for Heun’s method and the RK4 integrator, model predictions are needed to correct the predicted step or predict derivatives at interpolated timesteps. These additional model evaluations roughly double or quadruple the inference cost, which is consistent with the amount of extra derivatives that need to be calculated in the numerical schemes. Another interesting finding is that when the dynamics are too stiff or $\Delta t$ is too large, lower-order integrators can quickly become unstable, while state-prediction models can maintain an inaccurate state over time. As as whole, we find that neural surrogates can be faster than un-optimized numerical solvers, but no neural surrogate is currently faster or more accurate than a highly-optimized numerical solver. While this is rather discouraging for neural surrogates, it does reflect similar findings from other works [75, 26] and future work in adaptive step-sizing or novel architectures can help to further accelerate neural surrogates in this regime.
5.4 Limitations
As a final remark, we note some limitations of derivative prediction. Not all time-stepping schemes can be implemented with derivative prediction; in particular, it is not obvious how to extend this work to more sophisticated schemes such as implicit integrators. Furthermore, for steady-state prediction problems or boundary value problems, which could encompass phenomena such as Darcy flow or linear elasticity, derivative prediction would not work in its current formulation. Lastly, derivative prediction introduces temporal discretization constraints into model inference. This contradicts one of the main benefits of neural surrogates, which is that they accelerate simulation by taking large timesteps at the cost of convergence or accuracy guarantees. When considering this full picture, state and derivative prediction have their unique benefits and drawbacks, and the most suitable framework will depend heavily on the given problem and desired behavior of the neural surrogate. We hope that the experiments and discussion can help practitioners understand when and how to apply derivative and state prediction.
Model: FNO Unet Solver
Experiment Metric State Adams Heun RK4 State Adams Heun RK4 Full Half Quarter
NS Runtime (s) 0.163 0.158 0.301 0.622 0.359 0.359 0.691 1.379 3.508 2.692 2.549
Rollout Error 0.715 0.100 0.100 0.100 0.099 0.048 0.049 0.049 0.000 0.096 0.285
Kolm. Flow Runtime (s) 0.152 0.155 0.308 0.616 0.204 0.201 0.411 0.818 0.116 - -
Rollout Error 0.707 1e9 0.452 0.449 0.810 1e8 90.05 0.353 $0$
Table 5: Computational Cost. Comparison of computational cost of different models and a baseline solver on the Navier-Stokes and Kolmogorov Flow experiments. Runtimes are reported in seconds (s) for a full rollout, averaged for each sample in the validation set. Rollout errors are given as relative L2 error, with some models being unstable. Furthermore, solver baselines are coarsened to match neural surrogate accuracies when needed by solving at full ($n_{x}\times y$), half ($\frac{n_{x}}{2}\times\frac{n_{y}}{2}$), and quarter ($\frac{n_{x}}{4}\times\frac{n_{y}}{4}$) spatial resolutions.
@@ 6 Conclusion
We evaluate framing neural surrogates to predict temporal derivatives and use an ODE integrator during inference to produce solutions. Although this idea has been studied informally, this is the first work to broadly apply this framework and investigate its benefits and limitations. We find that derivative prediction can improve model accuracy and stability across architectures and PDE problems, and has useful data and inference flexibility. Furthermore, we compare this framework with other training modifications and study the effect of temporal resolution on model performance. Lastly, we hypothesize potential explanations for the improved stability and accuracy of derivative prediction and study the sources of error in derivative prediction. We hope that future work can apply derivative prediction across a wider variety problems and model architectures, and improve performance at coarser step sizes.